OpenStudy (anonymous):

If the circumference of the circle below is 25cm, what is the length of AB, to the nearest hunderedth?

OpenStudy (anonymous):

@Swageraid32 @cherio12 @mathstudent55 @zepdrix

OpenStudy (anonymous):

so first get the circumference

OpenStudy (anonymous):

@cherio12 2 * 3.14 * 25?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so, now the angle between a and b is 70 degrees. there is a total of 360 in a circle

OpenStudy (anonymous):

so you can do a ratio of 70/360 to get the fraction that a and b cover

OpenStudy (anonymous):

multiply that fraction by the circumference to get the length. Does that make sense?

OpenStudy (anonymous):

I got 157 as circumference :) I got 30.53 as my answer? @cherio12

OpenStudy (anonymous):

that's what i got

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem

OpenStudy (mathstudent55):

@needmathhelp7 @cherio12 The problem reads "the circumference of the circle below is 25 cm." You are given the circumference. All you need to do is the ratio of a 70 degree angle to the full 360 degrees of the circle. L = (k/360)(2)(pi)r L = (70/360)(25 cm) L = 4.86 cm

OpenStudy (anonymous):

bummers did i completely mess it up >.<

OpenStudy (anonymous):

yes...yes i did